# DACAD: Domain Adaptation Contrastive Learning for Anomaly Detection in Multivariate Time Series

### v2

@@ III-F Overall Loss in DACAD
III-G DACAD’s Inference
In the inference phase of the DACAD model, the primary objective is to identify anomalies in the target domain $T$. This is accomplished for each window $w^{t}$ in $T$. The anomaly detection is based on the concept of spatial separation in the feature space, as established during the training phase. The anomaly score for each window is derived from the spatial distance between the classifier’s output and a predefined centre $c$ in the feature space. This score is calculated as the squared Euclidean distance, as shown in Equation (6):
@@ III-G DACAD’s Inference
Where $\phi^{CL}(\phi^{R}(w^{t}))$ denotes the feature representation of the window $w^{t}$ after processing by the representation layer $\phi^{R}$ and the classifier layer $\phi^{CL}$. The distance measured here quantifies how much each window in the target domain deviates from the “normal” pattern, as defined by the centre $c$.
The anomaly score is a crucial metric in this process. A higher score suggests a significant deviation from the normative pattern, indicating a higher likelihood of the window being an anomaly. Conversely, a lower score implies that the window’s representation is closer to the centre, suggesting it is more likely to be normal.
In practical terms, the anomaly score can be thresholded to classify windows as either normal or anomalous. By effectively utilizing these anomaly scores, the DACAD model provides a robust mechanism for identifying anomalous patterns in unlabelled target domain data.
@@ IV Experiments
This section provides a comprehensive evaluation of DACAD, covering the experimental setups (Section V-A) and results (Section V-B and V-E), to clearly understand its performance and capabilities in different contexts.
@@ IV-A Datasets
We evaluate the proposed model and make comparisons across the four datasets, including the three most commonly used real benchmark datasets for TSAD and the dataset used for time series domain adaptation. The datasets are summarised in Table I.
@@
Boiler Fault Detection Dataset. [^3] https://github.com/DMIRLAB-Group/SASA-pytorch/tree/main/datasets/Boiler [8] The Boiler dataset includes sensor information from three separate boilers, with each boiler representing an individual domain. The objective of the learning process is to identify the malfunctioning blowdown valve in each boiler. Obtaining samples of faults is challenging due to their scarcity in the mechanical system. Therefore, it’s crucial to make use of both labeled source data and unlabeled target data to enhance the model’s ability to generalize.
@@ V Baselines
Below, we will provide an enhanced description of the UDA models for time series classification and anomaly detection. Additionally, we provide a description of the five prominent and state-of-the-art TSAD models that were used for comparison with DACAD. We have selected the models from different categories of TSAD, namely, unsupervised reconstruction-based (OmniAnomaly [44] and AnomalyTransformer [39]) model, unsupervised forecasting-based (THOC [47]), and self-supervised contrastive learning (TS2Vec [36] and DCdetector [53]) TSAD models.
@@
AE-LSTM [65] is an anomaly detection model using an LSTM-based encoder-decoder for multi-sensor data. It detects anomalies by reconstruction error of time series sequences.
RDC [^4] https://github.com/syorami/DDC-transfer-learning [66] is built upon deep domain confusion to enhance domain adaptation by maximizing domain invariance. It aligns feature distributions through a domain confusion objective, making learned representations indistinguishable across domains and improving predictive model performance on unseen domains.
VRADA [^5] https://github.com/floft/vrada [7] combines deep domain confusion [66] with variational recurrent adversarial deep domain adaptation [7] optimizing source domain label prediction, MMD, and domain discrimination using latent representations from the LSTM encoder. Concurrently, AE-LSTM’s reconstruction objective is utilized for anomaly detection.
@@
CLUDA [^7] https://github.com/oezyurty/CLUDA [9] is a novel framework developed for UDA in time series data and evaluated on time series classification. This framework utilizes a contrastive learning approach to learn domain-invariant semantics within multivariate time series data, aiming to preserve label information relevant to prediction tasks. CLUDA is distinctive as it is the first UDA framework designed to learn contextual representations of time series data, maintaining the integrity of label information. Its effectiveness is demonstrated through evaluations of various time series classification datasets.
@@
OmniAnomaly [^9] https://github.com/smallcowbaby/OmniAnomaly [44] is a model operating on an unsupervised basis, employing a Variational Autoencoder (VAE) to handle multivariate time series data. It identifies anomalies by evaluating the reconstruction likelihood of specific data windows.
@@
AnomalyTransformer [^12] https://github.com/thuml/Anomaly-Transformer [39] detects anomalies in time series data by using a novel ”Anomaly-Attention” mechanism to compute association discrepancies, enhancing the distinguishability between adjacent normal and abnormal points through a minimax strategy.
@@ V-A Experimental Setup
In our study, we evaluate several SOTA TSAD models including OmniAnomaly [44], TS2Vec [36], THOC [47], and DCdetector [53] and SOTA UDA models that support multivariate time series classification including VRADA [7] and CLUDA [9] on benchmark datasets previously mentioned in section I using their source code and best hyper-parameters as they stated to ensure a fair evaluation. The hyper-parameters used in our implementation are as follows: DACAD consists of a 3-layer TCN architecture with three different channel sizes [128, 256, 512] to capture temporal dependencies. The dimension of the representation is 1024. We use the same hyper-parameters across all datasets to evaluate DACAD: window size ($WS$) = 100, margin $m$ = 1, and we run our model for 20 epochs.
TABLE II: F1, AUPR, and AUROC results for various models on multivariate time series benchmark datasets (SMD, MSL, SMAP). The most optimal evaluation results are displayed in bold, while the second best ones are indicated by underline.
@@ V-B Baselines Comparison
Table II provides a comprehensive comparison of the performance of the different models on multivariate time series benchmark datasets. The performance measures include the F1 score, AUPR (Area Under the Precision-Recall Curve), and AUROC (Area Under the Receiver Operating Characteristic Curve). Despite Point Adjustment (PA) popularity in recent years, we do not use PA when calculating these metrics due to Siwon Kim’s findings  [67] that its application leads to an overestimation of a TSAD model’s capability and can bias results in favor of methods that produce extreme anomaly scores. Instead, we use conventional performance metrics for anomaly detection.
@@
TABLE III: Macro F1/AUROC results on SMD and Boiler dataset. The most optimal evaluation results are displayed in bold, while the second best ones are indicated by underline.
@@ V-C UDA Comparison
Table III, adapted from [13], now includes results of CLUDA and our model in the two last columns. This table provides a comprehensive comparison of the performances of various models on the SMD and Boiler datasets, as measured by Macro F1 and AUROC scores.
To assess our model against ContextDA [13] -the only exiting UDA for the TSAD model-, we use the same datasets and metrics reported in ContextDA’s main paper, as its code is unavailable. On the SMD dataset, ContextDA achieves an average macro F1 of 0.63 and AUROC of 0.75, whereas our model achieves a higher average macro F1 of 0.81 and AUROC of 0.86. Similarly, on the Boiler dataset, ContextDA’s average macro F1 is 0.50 and AUROC 0.65, compared to our model’s superior performance with an average macro F1 of 0.63 and AUROC of 0.71. In this comparative analysis, the DACAD model distinctly outperforms others, consistently achieving the highest scores in both Macro F1 and AUROC across the majority of test cases in both datasets. Its performance is not only superior but also remarkably stable across different scenarios within each dataset, showcasing its robustness and adaptability to diverse data challenges.
@@
In conclusion, while the DACAD model emerges as the most effective across most scenarios, the varying performances of different models across scenarios and datasets emphasise the need to carefully consider the unique characteristics and capabilities of each model when selecting the most appropriate one for a specific task. This nuanced approach is crucial in leveraging the strengths of each model to achieve optimal results.
@@ V-D Scalability Test
Scalability tests are crucial to evaluate a model’s performance as data scales up in size and complexity. The plots in Fig.2 provide a comprehensive comparison of DACAD’s scalability with other models described in section V-B. The experiment measures execution time and peak memory usage in relation to the length of time series and their dimensionality. For the dimensionality test, eight synthetic time series datasets with varying dimensions (ranging from 8 to 1024, doubling each step) and a fixed length of 2,000 are generated using the NeurIPS-TS [69]. Another set of eight synthetic datasets with a fixed dimension of 8 varying lengths (ranging from 2,000 to 256,000) are created for the scale-up test concerning length. We report the execution time and memory usage of all models with the same parameter settings as in section V-B.
@@ V-E Ablation Study
Table IV offers several key insights into the impact of different loss components on the MSL dataset using F-5 as a source. Removing the target self-supervised CL (w/o $\mathcal{L}_{\text{SelfCont}}$) leads to lower metrics (F1: 0.481, AUPR: 0.495, AUROC: 0.697). Moreover, excluding source-supervised CL (w/o $\mathcal{L}_{\text{SupCont}}$) reduces effectiveness (F1: 0.463, AUPR: 0.427, AUROC: 0.639), highlighting its role in capturing source-specific features which are crucial for the model’s overall accuracy. Similarly, omitting the discriminator component results in performance reduction (F1: 0.471, AUPR: 0.484, AUROC: 0.699). However, the most significant decline occurs without the classifier (F1: 0.299, AUPR: 0.170, AUROC: 0.503), underscoring its crucial role in effectively distinguishing between normal/anomaly classes. Overall, the best results (F1: 0.595, AUPR: 0.554, AUROC: 0.787) are achieved with all components, highlighting the effectiveness of an integrated approach.
@@
To elucidate the effectiveness of UDA within DACAD, we examine the feature representations from the MSL dataset, as illustrated in Fig. 3. It presents the t-SNE 2D embeddings of DACAD feature representations $\phi^{CL}(\phi^{R}(w))$ for the MSL dataset. Each point represents a time series window, which can be normal, anomalous, or anomaly-injected. These representations highlight the domain discrepancies between source and target entities and demonstrate how DACAD aligns the time series window features effectively. The comparison of feature representations with and without UDA reveals a significant domain shift when UDA is not employed, between source (entity F-5) and the target (entity T-5) within the MSL dataset.
@@
Effect of anomaly injection: We study the impact of anomaly injection in Table VI, on the MSL dataset when using F-5 as a source. It shows that anomaly injection significantly improves all metrics (F1: 0.595, AUPR: 0.554, AUROC: 0.787), enhancing the model’s ability to differentiate between normal and anomalous patterns, thereby improving DACAD’s overall accuracy. Without anomaly injection, there is a notable decline in performance, emphasizing its role in precision. The higher standard deviation in AUROC scores without injection suggests more variability and less stability in the model performance. The study underscores the vital role of anomaly injection in improving anomaly detection models. It reveals that incorporating anomaly injection not only boosts detection accuracy but also enhances the model’s overall stability and reliability.
@@ VI Conclusion
The DACAD model is an innovative approach to TSAD that is particularly effective in environments with limited labeled data. By combining domain adaptation and contrastive learning, it applies labelled anomaly data from one domain to detect anomalies in another. Its anomaly injection mechanism, introducing a spectrum of synthetic anomalies, significantly bolsters the model’s adaptability and robustness across various domains. Our evaluations on diverse real-world datasets establish DACAD’s superiority in handling domain shifts and outperforming existing models. Its capability to generalize and accurately detect anomalies, regardless of the scarcity of labeled data in the target domain, represents a significant contribution to TSAD. In future work, we aim to refine the anomaly injection process further, enhancing the model’s ability to simulate a broader range of anomalous patterns. Additionally, we plan to evolve the model to encompass univariate time series analysis, broadening its scope and utility.

### v3

@@ III-F Overall Loss in DACAD
III-G Role of Anomaly Injection in Convergence
Under standard assumptions in nonconvex optimization, the anomaly injection mechanism in DACAD enforces a minimum margin $m$ between normal and anomalous embeddings. This guarantees that the contrastive losses remain strictly positive unless the model achieves the desired separation, thereby preventing degenerate solutions such as feature collapse. Assuming that the feature extractor is $L$-Lipschitz and that the loss functions are smooth with Lipschitz-continuous gradients [59], an update via stochastic gradient descent with a sufficiently small learning rate ensures that the overall loss is lower bounded. Moreover, under these conditions, the loss locally satisfies a Polyak–Łojasiewicz condition, which implies that the expected gradient norm decreases at a rate of $\mathcal{O}(1/K)$ over $K$ iterations [60]. This result formally guarantees convergence to a stationary point.
III-H DACAD’s Inference
In the inference phase of the DACAD model, the primary objective is to identify anomalies in the target domain $T$. This is accomplished for each window $w^{t}$ in $T$. The anomaly detection is based on the concept of spatial separation in the feature space, as established during the training phase. The anomaly score for each window is derived from the spatial distance between the classifier’s output and a learnable center $c$ in the feature space. This score is calculated as the squared Euclidean distance, as shown in Equation  (6):
@@ III-H DACAD’s Inference
Where $\phi^{CL}(\phi^{R}(w^{t}))$ denotes the feature representation of the window $w^{t}$ after processing by the representation layer $\phi^{R}$ and the classifier layer $\phi^{CL}$. The distance measured here quantifies how much each window in the target domain deviates from the “normal” pattern, as defined by the center $c$.
The anomaly score is a crucial metric in this process. A higher score suggests a significant deviation from the normative pattern, indicating a higher likelihood of the window being an anomaly. Conversely, a lower score implies that the window’s representation is closer to the center, suggesting it is more likely to be normal. In practical terms, the anomaly score can be thresholded to classify windows as either normal or anomalous. By effectively utilizing these anomaly scores, the DACAD model provides a robust mechanism for identifying anomalous patterns in unlabeled target domain data.
@@ IV Experiments
This section provides a comprehensive evaluation of DACAD, covering the experimental setups (Section V-A) and results (Section V-B and V-E) to clearly understand its performance and capabilities in different contexts.
@@ IV-A Datasets
We evaluate the proposed model and make comparisons across the four datasets, including the three most commonly used real benchmark datasets for TSAD and the dataset used for time series domain adaptation. The datasets are summarized in Table I. Since DACAD utilizes UDA, it is designed to work in scenarios where datasets contain multiple entities. Domain adaptation relies on inter-domain feature alignment, requiring at least one entity as the source and one as the target.
@@
Boiler Fault Detection Dataset. [^4] https://github.com/DMIRLAB-Group/SASA-pytorch/tree/main/datasets/Boiler [8] The Boiler dataset includes sensor information from three separate boilers, with each boiler representing an individual domain. The objective of the learning process is to identify the malfunctioning blowdown valve in each boiler. Obtaining samples of faults is challenging due to their scarcity in the mechanical system.
@@ V Baselines
Below, we will provide an enhanced description of the UDA models for time series classification and anomaly detection. Additionally, we provide a description of the five prominent and state-of-the-art TSAD models that were used for comparison with DACAD. We have selected the models from different categories of TSAD, namely, unsupervised reconstruction-based (OmniAnomaly [45] and AnomalyTransformer [41]) models, unsupervised forecasting-based (THOC [47]), and self-supervised contrastive learning (TS2Vec [38] and DCdetector [52]) TSAD models.
@@
AE-LSTM [62] is an anomaly detection model using an LSTM-based encoder-decoder for multi-sensor data. It detects anomalies by the reconstruction error of time series sequences.
RDC [^5] https://github.com/syorami/DDC-transfer-learning [63] is built upon deep domain confusion to enhance domain adaptation by maximizing domain invariance. It aligns feature distributions through a domain confusion objective, making learned representations indistinguishable across domains and improving model performance on unseen domains.
VRADA [^6] https://github.com/floft/vrada [7] combines deep domain confusion [63] with variational recurrent adversarial deep domain adaptation [7], optimizing source domain label prediction, MMD, and domain discrimination using latent representations from the LSTM encoder. Concurrently, AE-LSTM’s reconstruction objective is utilized for anomaly detection.
@@
CLUDA [^8] https://github.com/oezyurty/CLUDA [9] is a novel UDA framework for time series data, leveraging contrastive learning to capture domain-invariant semantics while preserving label information. It is the first UDA framework designed for contextual representation learning in MTS, with effectiveness demonstrated on multiple time series classification datasets.
@@
OmniAnomaly [^10] https://github.com/smallcowbaby/OmniAnomaly [45] is a model operating on an unsupervised basis, employing a Variational Autoencoder (VAE) to handle MTS data. It identifies anomalies by evaluating the reconstruction likelihood of specific data windows.
@@
AnomalyTransformer [^13] https://github.com/thuml/Anomaly-Transformer [41] detects anomalies in time series data by using a novel “Anomaly-Attention” mechanism to compute association discrepancies, enhancing the distinguishability between adjacent normal and abnormal points through a minimax strategy.
@@ V-A Experimental Setup
In our study, we evaluate several SOTA TSAD models, including OmniAnomaly [45], TS2Vec [38], THOC [47], and DCdetector [52] and SOTA UDA models that support MTS classification including VRADA [7] and CLUDA [9] on benchmark datasets previously mentioned in section I using their source code and best hyper-parameters as they stated to ensure a fair evaluation. The hyperparameters used in our implementation are as follows: DACAD consists of a 3-layer TCN architecture with three different channel sizes [128, 256, 512] to capture temporal dependencies. The dimension of the representation is 1024. We use the same hyperparameters across all datasets to evaluate DACAD: window size ($W_{S}$) = 100, margin $m$ = 1, and we run our model for 20 epochs.
TABLE II: F1, AUPR, and AUROC results for various models on multivariate time series benchmark datasets (SMD, MSL, SMAP). The most optimal evaluation results are displayed in bold, while the second-best ones are indicated by underline.
@@ V-B Baselines Comparison
Table II provides a comprehensive comparison of the performance of the different models on MTS benchmark datasets. The performance measurements include the F1 score (best F1 score), AUPR (Area Under the Precision-Recall Curve), and AUROC (Area Under the Receiver Operating Characteristic Curve). Despite Point Adjustment (PA) popularity in recent years, we do not use PA when calculating these metrics due to Siwon Kim’s findings  [64] that its application leads to an overestimation of a TSAD model’s capability and can bias results in favor of methods that produce extreme anomaly scores. Instead, we use conventional performance metrics for anomaly detection.
@@
TABLE III: Macro F1/AUROC results on SMD and Boiler dataset. The most optimal evaluation results are displayed in bold, while the second-best ones are indicated by underline.
@@ V-C UDA Comparison
Table III, adapted from [14], now includes results of CLUDA and our model in the last two columns. This table provides a comprehensive comparison of the performances of various models on the SMD and Boiler datasets, as measured by Macro F1 and AUROC scores.
To assess our model against ContextDA [14] — the only existing UDA for the TSAD model — we use the same datasets and metrics reported in ContextDA’s main paper, as its code is unavailable. On the SMD dataset, ContextDA achieves an average macro F1 of 0.63 and AUROC of 0.75, whereas our model achieves a higher average macro F1 of 0.81 and AUROC of 0.86. Similarly, on the Boiler dataset, ContextDA’s average macro F1 is 0.50 and AUROC 0.65, compared to our model’s superior performance with an average macro F1 of 0.63 and AUROC of 0.71. In this comparative analysis, the DACAD model distinctly outperforms others, consistently achieving the highest scores in both Macro F1 and AUROC across the majority of test cases in both datasets. Its performance is not only superior but also remarkably stable across different scenarios within each dataset, showcasing its robustness and adaptability to diverse data challenges.
@@
In conclusion, while the DACAD model emerges as the most effective across most scenarios, the varying performances of different models across scenarios and datasets highlight the importance of carefully evaluating each model’s unique characteristics and capabilities. A thoughtful selection process is essential to ensuring the most suitable model is chosen for a given task. This subtle approach is crucial in leveraging each model’s strengths and achieving optimal results.
@@ V-D Scalability Test
Scalability tests are crucial to evaluate a model’s performance as data scales up in size and complexity. The plots in Fig. 2 provide a comprehensive comparison of DACAD’s scalability with other models described in section V-B. The experiment measures execution time and peak memory usage in relation to the length of the time series and their dimensionality. For the dimensionality test, eight synthetic time series datasets with varying dimensions (ranging from 8 to 1024, doubling each step) and a fixed length of 2,000 are generated using the NeurIPS-TS [66]. Another set of eight synthetic datasets with a fixed dimension of 8 varying lengths (ranging from 2,000 to 256,000) is created for the scale-up test concerning length. We report the execution time and memory usage of all models with the same parameter settings as in section V-B.
@@ V-E Ablation Study
Table IV offers several key insights into the impact of different loss components on the MSL dataset using F-5 as a source. Removing the target self-supervised CL (w/o $\mathcal{L}_{\text{SelfCont}}$) leads to lower metrics (F1: 0.481, AUPR: 0.495, AUROC: 0.697). Moreover, excluding source-supervised CL (w/o $\mathcal{L}_{\text{SupCont}}$) reduces effectiveness (F1: 0.463, AUPR: 0.427, AUROC: 0.639), highlighting its role in capturing source-specific features, which are crucial for the model’s overall accuracy. Similarly, omitting the discriminator component results in performance reduction (F1: 0.471, AUPR: 0.484, AUROC: 0.699). However, the most significant decline occurs without the classifier (F1: 0.299, AUPR: 0.170, AUROC: 0.503), underscoring its crucial role in effectively distinguishing between normal/anomaly classes. Overall, the best results (F1: 0.595, AUPR: 0.554, AUROC: 0.787) are achieved with all components, highlighting the effectiveness of an integrated approach.
@@
To elucidate the effectiveness of UDA within DACAD, we examine the feature representations from the MSL dataset, as illustrated in Fig. 3. It presents the t-SNE 2D embeddings of DACAD feature representations $\phi^{CL}(\phi^{R}(w))$ for the MSL dataset. Each point represents a time series window, which can be normal, anomalous, or anomaly-injected. These representations highlight the domain discrepancies between source and target entities and demonstrate how DACAD aligns the time series window features effectively. The comparison of feature representations with and without UDA reveals a significant domain shift when UDA is not employed between the source (entity F-5) and the target (entity T-5) within the MSL dataset.
@@
Effect of anomaly injection: We study the impact of anomaly injection in Table VI on the MSL dataset when using F-5 as a source. It shows that anomaly injection significantly improves all metrics (F1: 0.595, AUPR: 0.554, AUROC: 0.787), enhancing the model’s ability to differentiate between normal and anomalous patterns, thereby improving DACAD’s overall accuracy. Without anomaly injection, there is a notable decline in performance, emphasizing its role in precision. The higher standard deviation in AUROC scores without injection suggests more variability and less stability in the model performance. The study underscores the vital role of anomaly injection in improving anomaly detection models. It reveals that incorporating anomaly injection not only boosts detection accuracy but also enhances the model’s overall stability.
@@ VI Conclusion
The DACAD model is an innovative approach to TSAD that is particularly effective in environments with limited labeled data. By combining domain adaptation and contrastive learning, it applies labeled anomaly data from one domain to detect anomalies in another. Its anomaly injection mechanism, introducing a spectrum of synthetic anomalies, significantly bolsters the model’s adaptability and robustness across various domains. Our evaluations on diverse real-world datasets establish DACAD’s superiority in handling domain shifts and outperforming existing models. Its capability to generalize and accurately detect anomalies, regardless of the scarcity of labeled data in the target domain, represents a significant contribution to TSAD. In future work, we aim to refine the anomaly injection process further, enhancing the model’s ability to simulate a broader range of anomalous patterns. Additionally, we plan to evolve the model to encompass univariate time series analysis, broadening its scope and utility. Furthermore, we intend to explore alternative feature extraction architectures beyond TCN to assess their impact on model performance.